Fermi National Laboratory
 

Operation Summary for October 4, 2010 - October 6, 2010

 

Monday Oct. 4

The day shift began with Operations monitoring the Tevatron (TeV) store 8141 (101.22E30) and the Recycler stash (206.1E10), with the Antiproton Source (Pbar) stacking (22.6E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with NuMI off for target work, with MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8141 at 2:40 PM and put the TeV into HEP shot setup. Switchyard (SY) personnel conducted studies.

During the evening shift, Operations established store 8142 at 4:15 PM with an average initial luminosity (AIL) of 300E30. Operations reported, at 10 PM, that the Meson Septa (MSEPS) developed a vacuum leak. Operations called in Mechanical Support personnel to investigate.

Tuesday Oct. 5

The midnight shift began with Operations monitoring store 8142 (92.08E30) and the stash (242.4E10), with Pbar stacking (11.9E10), and with MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8142 at 3:45 AM. Operations prepared enclosure B for access so that Mechanical Support personnel could install vacuum pumps on MSEPS. Operations put the TeV into shot setup and established store 8143 at 7:10 AM with an AIL of 306E30.

During the day shift, at 8:30 AM, Operations aborted store 8143 when the Cryo System Coordinator deemed it necessary to repair a high-pressure helium leak at TeV sector C4. This triggered accesses into Enclosure B to further investigate the MSEPS vacuum problem and accesses to the following enclosures for routine maintenance: Tevatron (including the collision halls), Pbar (Rings, Transport, Pre-Vault, and Pre-Target), MI-12B, H- Source, and Booster. Operations reported that Enclosure B and Transfer Hall accesses were still underway at the end of the shift. Operations delivered minimal beam to users and the Recycler maintained a stash of ~85E10.

During the evening shift, Operations established beam to Pbar and MiniBooNE at 4:15 PM. Everyone was out of the Transfer Hall and Enclosure B by 6:30 PM. Operations reported a TeV sector C2 dump failure during turn-on and called in EE Support personnel to investigate. Operators completed the TeV check out and turn-on at 11:30 PM. Operations put the TeV into shot setup 11:45 PM.

Wednesday Oct. 6

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (5.2E10), with Pbar stacking (29.5E10), and with MiniBooNE taking beam.

During the midnight shift, Operations established store 8145 at 12:45 AM with an AIL of 238E30. Operations reported an issue with an ARCnet controller at AP30 that required attention from Controls personnel.

The Plans for Wednesday

The plans for today are to terminate store 8145 and then establish a new store, so that MI and Vacuum personnel can access MI to investigate a Lambertson vacuum issue.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 10/06/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory